# Large Language Models Assume People are More Rational than We Really are

### v3

@@ 1 Introduction
Figure 1: Two tasks we use to assess the implicit assumptions that LLMs make about human decision-making. (A) Predicting choices between gambles. Each gamble is described by the probabilities and values of different outcomes, and the goal is to predict which gamble people will choose. (B) Inferring preferences from choices. Here, a person chooses one of many sets of objects and the goal is to infer their preferences based on that choice.
@@
Every day, our actions are based on the countless decisions that we make reflecting our internal goals and beliefs about the world. Through our interactions with others, we are able to effortlessly predict how other people would act from their goals and beliefs, and infer others’ goals and beliefs when observing their choices. These abilities — termed forward- and inverse-modeling in cognitive science (Ho and Griffiths, 2021) — are characteristic of the implicit decision-making models that we form of others, and are crucial to interpersonal communication and learning (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020). However, while these abilities are inherent in people, the accuracy and consistency of decision-making models in Large Language Models (LLMs) is unknown. As LLMs become widely used as the basis for building AI agents that interact with or simulate people, it is important to ask what decision-making models LLMs implicitly use: the ability to predict and interpret people’s behavior is a precursor to identifying effective ways to provide assistance, simulating the helpfulness or harmlessness of a response, and learning individuals’ values and preferences, all of which are principal to the development of safe and beneficial AI systems.
@@
Though LLMs have become increasingly capable of conducting reasoning and conversing with humans, there is no guarantee that their implicit representations of humans align with how we actually behave. Methods such as Proximal Policy Optimization (Schulman et al., 2017) and Direct Preference Optimization (Rafailov et al., 2023) can be used to tune models on explicitly declared human preferences (Ouyang et al., 2022), but training data such as blog posts, news articles, and books go through rounds of editing that remove logical fallacies and other mistakes, leading to more “perfect” content being used for training (Cui et al., 2023; Zhou et al., 2024a). While this improves the quality of generation, it may also lead LLMs to develop mistaken impressions of how humans actually behave. Furthermore, many existing measures comparing LLM simulations to human behavior use human perception to measure similarity (Park et al., 2023; Jones and Bergen, 2023; Jakesch et al., 2023; Hämäläinen et al., 2023), but existing psychology literature has shown that people’s own perceptions of others may be more rational than they actually are (Jern et al., 2017). Thus, an AI system that appears human-like to the naked eye may not actually correspond to human behavior.
@@
This phenomenon is not without precedent — early economists did not view the assumption of human rationality as problematic (Smith, 1776; Mill, 1836) and built sophisticated models and policies (Downs, 1957; Coleman, 1994; Schelling, 1980; Dunleavy, 2014, inter alia), until psychologists showed just how systematic and widespread its failures were in accounting for human behavior (Tversky and Kahneman, 1974; Kahneman and Tversky, 1979). As LLM-powered systems become more widely-used, misaligned representations of humans — which can lead to mismatched beliefs and failure to follow instructions (Milli and Dragan, 2020) — may pose a greater toll on various downstream applications. But how can we study these implicit decision-making models without relying on human judgments?
@@
To explore the implicit decision-making models of LLMs, we leverage two existing experimental paradigms from psychology — a risky choice task where humans choose between gambles (Bourgin et al., 2019, Figure 1A), and an inference task where people reason about others’ subjective utilities after observing their decisions (Jern et al., 2017, Figure 1B). In the former, we compare choices that participants made with LLM simulations and predictions of those choices (forward modeling; Ho and Griffiths, 2021), while in the latter we compare the inferences drawn between models and humans over the same observations (inverse modeling). These two paradigms are connected by the same foundational theoretical models of human decision-making, under which people develop mental models of others’ goals, utilities, and decision spaces, and use these to 1) construct predictions about how others will behave given their beliefs, and 2) infer what people believe based on their decisions (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021).
@@
Through these experiments, we find that LLMs model people as highly rational decision makers. For forward modeling, we discover that when prompting with chain-of-thought (CoT; Wei et al., 2022), LLMs consistently predict that humans act more rationally than they do; today’s most capable LLMs produced correlations greater than 0.94 with the rational model of choosing based on the maximum expected value, but humans only have a correlation of 0.48. Additionally, asking LLMs to simulate the decision with CoT also yields highly rational outcomes that are only moderately correlated with human behavior. In both cases, zero-shot prompting generates noisy outcomes that are only moderately correlated with either rational models or humans.
@@
In the inverse modeling paradigm, we find that the inferences that LLMs make from peoples’ choices are also consistent with the assumption that humans are rational actors. Across two contexts, LLMs’ inferences positively correlate with predictions from rational models, with these correlations steadily increasing with model capabilities and from zero-shot to CoT (0.20 with Llama-3-8B zero-shot; 0.95 with GPT-4o CoT). However, the psychology literature suggests that despite deviating from rationality in their own choices, people’s inferences from the behavior of others are consistent with assuming humans are rational agents (Baker et al., 2009; Lucas et al., 2014; Jern et al., 2017; Jara-Ettinger et al., 2020). Consequently, we also find that the inferences made by LLMs from others’ choices are very highly correlated with the same inferences made by people. Thus, while LLMs are not accurate at simulating or predicting human behavior, LLMs’ assumption that people are more rational than we really are aligns with the assumption that people make when interpreting one another’s behavior.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce related work across LLMs, alignment, and models of decision-making in cognitive science. Section 3 and Section 4 describe the forward and inverse modeling experiments and results, detailing our analyses and the correlations between LLMs, humans, and existing rational theories. Finally, in Section 5, we discuss potential insights related to our work, including implications towards the feasibility of simulating humans using LLMs, how RLHF can bias models towards aligning with what humans do instead of what humans expect others to do (and vice versa), and the potential of using cognitive science to inspire different forms of alignment research.
@@ 2.1 Aligning large language models with humans
Large Language Models are typically aligned with human preferences through Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022). Training with human preference data has been shown to enhance reasoning in LLMs  (Havrilla et al., 2024). The impressive capacities of the resulting models (e.g., Bubeck et al., 2023) have sparked interest across various fields in using them to model (e.g., Binz and Schulz, 2023a; Macmillan-Scott and Musolesi, 2024) and simulate (e.g., Park et al., 2022; Argyle et al., 2023; Liu et al., 2023) human behavior. Even though aligned LLMs can effectively approximate average human judgments (Rathje et al., 2023), these models still exhibit biases and output hallucinations (Jiang et al., 2024; Bai et al., 2024; Anwar et al., 2024). Moreover, LLMs may not be adept at capturing trade-offs in human behavior (Liu et al., 2024; Coletta et al., 2024) or situations with information asymmetry (Zhou et al., 2024b).
Methods from cognitive science are increasingly being used to study LLMs (Hardy et al., 2023; Coda-Forno et al., 2024; Binz and Schulz, 2023b). A particularly contentious debate is whether LLMs exhibit Theory of Mind, the ability to model others’ mental states which may be different than their own (Premack and Woodruff, 1978). Several studies have shown evidence for (Kosinski, 2024) and against (Sap et al., 2022; Ullman, 2023) this claim, and have explored the ability of LLMs to accurately emulate various personas that should result in changes in their decision-making behavior (Salewski et al., 2024). Our analysis provides a quantitative approach to engaging with this debate, as inverse decision-making is a specific form of Theory of Mind.
@@ 2.2 Forward and inverse models of human decision-making
One of the most basic and extensively-studied problems in decision-making is the risky choice task (Kahneman and Tversky, 1979; Edwards, 1954; Peterson et al., 2021; Bourgin et al., 2019), where people choose among gambles with different outcome probabilities and payoffs. The rational action is to choose the gamble with the highest expected value, calculated by summing the product of the probabilies and the values of the outcomes. Humans have been described as deviating from rationality in a fourfold pattern: risk seeking for small probability gains, risk averse for small probability losses, risk averse for moderate and large probability gains, and risk seeking for moderate and large probability losses (Tversky and Kahneman, 1989; Kahneman and Tversky, 1979). Peterson et al. (2021) collected the choices13k dataset, a large dataset of over 13,000 human choices between different gambles, and showed that people’s decisions in this setting could be captured by relatively simple machine learning models. Binz and Schulz (2023a) used these data to fine-tune an LLM, achieving similar performance. Chen et al. (2023) built a risky choice dataset and found that GPT-3.5 makes economically rational decisions, which we replicate in task 3 of our forward modeling experiments on the choices13k dataset across various LLMs.
People are able to infer an agent’s beliefs, desires, and percepts from their actions (Baker et al., 2017; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021). These inferences are typically modeled by assuming that people employ a forward model of decision-making — typically a noisy rational model — and use Bayes’ rule to invert that model (for more details, see Section 4). The study that we focus on, by Jern et al. (2017), was intended to directly test this assumption. Similar approaches have been used to improve human-AI interaction (e.g., Dragan et al., 2013; Sadigh et al., 2016). Recent work has extended these models using LLMs to make inferences from utterances as well as actions (Zhi-Xuan et al., 2024; Ying et al., 2023).
@@ 3 Forward modeling: Predicting which gamble people will choose
Our forward modeling experiments used a traditional risky choice task from psychology, where participants choose between two options that differ in the probability of returning rewards.
@@
To comprehensively understand how LLMs empirically capture human intent and align with actual human decisions, we designed three forward modeling paradigms. First, we asked LLMs to predict the decisions that a human participant would make. Second, we asked LLMs to predict the proportion of participants that would select each option. Third, we instructed LLMs to simulate participants by making the decisions themselves. For each task we implemented both zero-shot and chain-of-thought (CoT; Wei et al., 2022) prompting. Our code is available at https://github.com/theryanl/LLM-rationality.
@@
Human choice data came from the choices13k dataset (Peterson et al., 2021; Bourgin et al., 2019), a comprehensive collection of 13,006 risky choice problems. Each choice problem consists of two options, where each option is a set of rewards with corresponding probabilities (e.g., would you rather win \$5 with probability 1 or take a 0.5 probability of winning \$7?). The dataset includes the proportion of people who selected each option for each problem. Our analyses used a subset of 9,831 of these problems that were not “ambiguous” (where probabilities were not shown) or lacked feedback. We used the data to evaluate the alignment of LLMs in each of the three paradigms with actual human choices.
@@
We evaluated the following general-purpose models, including both open-sourced and closed-sourced models: Llama-3-8B, Llama-3-70B, Claude 3 Opus, GPT-4-Turbo (0125-preview), and GPT-4o. We implemented experiments on the full choices13k dataset for zero-shot and chain-of-thought (CoT) prompting across the three different tasks mentioned above for all models besides Claude. For Claude 3 Opus, we only evaluated the first task — predicting what a human participant’s choice might be — due to cost limitations.
For the experiments predicting and simulating individual human decisions (tasks 1 & 3) and models Llama-3-8B, Llama-3-70B, and GPT-4-Turbo, we conducted zero-shot experiments with $n$ completions, where $n$ is the number of participants that made the same decision in the choices13k dataset (ranging from 15 to 33). Each completion involved predicting or simulating a single human participant’s decision, with temperature set to 0.7 to maintain sample diversity. For the corresponding CoT experiments, we observed that responses were completely deterministic at temperature 0.7 on a random subset of 1000 decisions, and thus prompted for 1 completion with temperature 0.0 across the entire dataset. For the same reason, we ran the experiments predicting the proportion of humans who would select an option (task 2) for 1 completion for both zero-shot prompting and CoT prompting with temperature 0.0.
We adapted a prompt template previously used by Binz and Schulz (2023a, see Appendix A). For each choice problem, we first introduced the decision context of the choices13k dataset (i.e., the idea of choosing between gambling options) before providing each option’s probabilities and associated rewards in dollars. Finally, we asked what the participant(s) would choose (for predicting decisions) or what “you” would choose (when acting as a human participant). To minimize any positional biases (Wang et al., 2023), we shuffled the order of the options presented in the prompt. Example prompts are in Appendix A.

### v4

@@ 1 Introduction
Every day, our actions are based on decisions that reflect our internal goals and beliefs about the world. Through countless interactions with others, we are able to effortlessly predict how other people would act from their goals and beliefs, and infer others’ goals and beliefs when observing their actions. These abilities — termed forward- and inverse-modeling in cognitive science (Ho and Griffiths, 2021) — are characteristic of the implicit mental decision-making models that we form of others, and are crucial to interpersonal communication and learning (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020). However, while these abilities are inherent in people, the consistency and accuracy of decision-making models in Large Language Models (LLMs) is unknown. As LLMs become widely used as the basis for building AI agents that interact with or even simulate people, it is important to ask what decision-making models LLMs implicitly use: the ability to predict and interpret people’s behavior is a precursor to identifying effective ways to provide assistance, simulating the helpfulness or harmlessness of a response, and learning individuals’ values and preferences, all of which are principal to the development of safe and beneficial AI systems.
@@
Though LLMs have become increasingly capable of conducting reasoning and conversing with humans, there is no guarantee that their implicit representations of humans align with how we behave. Methods such as Proximal Policy Optimization (Schulman et al., 2017) and Direct Preference Optimization (Rafailov et al., 2023) can be used to tune models on explicitly declared human preferences (Ouyang et al., 2022), but training data such as blog posts, news articles, and books go through rounds of editing that remove logical fallacies and mistakes, leading to more “perfect” content being used for training (Cui et al., 2023; Zhou et al., 2024a). While this improves the quality of generation, it may also lead LLMs to develop mistaken impressions of how humans actually behave. In addition, many of our methods comparing LLMs to human behavior rely on human perception to measure similarity (Park et al., 2023; Jones and Bergen, 2023; Jakesch et al., 2023; Hämäläinen et al., 2023), but the psychology literature has shown that people’s perceptions of other people are flawed — we expect others to be more rational than they are (e.g., Jern et al., 2017). Thus, AI systems that act rationally can appear human-like to the naked eye, while not actually capturing human behavior.
Figure 1: Two tasks we use to assess the implicit assumptions that LLMs make about human decision-making. (A) Predicting choices between gambles. Each gamble is described by the probabilities and values of different outcomes, and the goal is to predict what people will choose.
(B) Inferring preferences from choices. Here, a person chooses one of many sets of objects and the goal is to infer their preferences based on that choice.
@@
This phenomenon is not without precedent — early economists embraced the assumption of human rationality (Smith, 1776; Mill, 1836) and built sophisticated models and policies around it (Downs, 1957; Coleman, 1994; Schelling, 1980; Dunleavy, 2014, inter alia), before psychologists showed just how systematic and widespread its failures were in accounting for human behavior (Tversky and Kahneman, 1974; Kahneman and Tversky, 1979). As LLM-powered systems become more widely-used, misaligned representations of humans — which can lead to mismatched beliefs and failure to follow instructions (Milli and Dragan, 2020) — pose a toll on downstream applications. But how can we study LLMs’ implicit decision-making models without being affected by the human bias to assume others act rationally?
@@
To explore the implicit decision-making models of LLMs, we leverage two experimental paradigms from psychology — a risky choice task where people choose between gambles (Kahneman and Tversky, 1979, Figure 1A), and an inference task where people infer others’ subjective utilities after observing their decisions (Jern et al., 2017, Figure 1B). In the former, we compare choices that participants made with LLM simulations and predictions across a large dataset of over 13,000 risky decisions (Bourgin et al., 2019), while in the latter we compare the inferences drawn by LLMs and by humans over a set of 47 observations (Jern et al., 2017). These two paradigms are connected by foundational theoretical models of human decision-making, under which people develop mental models of others’ goals, utilities, and decisions, and use them to 1) construct predictions about how others will behave given their beliefs, and 2) infer what people believe based on their decisions (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021).
@@
Through our experiments, we find that LLMs model people as highly rational decision makers. In the forward modeling task, when prompting with chain-of-thought (CoT; Wei et al., 2022), LLMs consistently predict that people act more rationally than we do. For example, GPT-4o produced a Spearman correlation of $0.94$ with the rational model of choosing the maximum expected value, but humans only have a correlation of $0.48$. Asking LLMs to simulate the decision with CoT also yields highly rational outcomes ($0.90$ for GPT-4o) that are only moderately correlated with human behavior. In both cases, zero-shot prompting generates noisy outcomes only somewhat correlated with either rational models or humans.
@@
In the inverse modeling task, inferences that LLMs made from peoples’ choices were also consistent with assuming that humans are rational. Across two contexts, LLMs’ inferences positively correlated with predictions from rational models, where correlations increased with model capability and reasoning (0.20 for Llama-3-8B zero-shot; 0.95 for GPT-4o CoT). Interestingly, psychology literature finds that people, despite deviating from rationality in their own choices, make inferences from the behavior of others as if they were rational (Baker et al., 2009; Lucas et al., 2014; Jern et al., 2017; Jara-Ettinger et al., 2020). Thus, the inferences drawn from others’ decisions by people and LLMs should actually be similar. We find support for this: Inferences made by LLMs are highly correlated with the same inferences made by people ($\rho=0.97$ with GPT-4o CoT). Thus, while LLMs are not accurate at simulating or predicting human behavior, LLMs’ incorrect assumption that people are more rational aligns with the assumption that people also make when interpreting another’s behavior.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce related work across LLMs, alignment, and decision-making in cognitive science. Section 3 and Section 4 describe the forward and inverse modeling experiments and results, detailing our analyses and the relationships between LLMs, humans, and existing rational theories. Finally, in Section 5 we discuss insights including new challenges for simulating humans using LLMs, misspecification of “alignment” in decision-making, and potential implications for training LLMs using different paradigms.
@@ 2.1 Aligning large language models with humans
Large Language Models are typically aligned with human preferences through Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022). Training with human preference data has been shown to enhance reasoning (Havrilla et al., 2024), and the impressive capacities of resulting models (e.g., Bubeck et al., 2023) have sparked interest across various fields in using them to model (e.g., Binz and Schulz, 2023a; Macmillan-Scott and Musolesi, 2024) and simulate (e.g., Park et al., 2022; Argyle et al., 2023; Liu et al., 2023; Salewski et al., 2024) human behavior. However, these models still exhibit biases and hallucinations (e.g., Jiang et al., 2024; Bai et al., 2024; Anwar et al., 2024), and may not be adept at capturing trade-offs in human behavior (Liu et al., 2024b; Coletta et al., 2024) or situations with information asymmetry (Zhou et al., 2024b).
Methods from cognitive science are increasingly being used to study LLMs (Coda-Forno et al., 2024; Liu et al., 2024b; Binz and Schulz, 2023b; Liu et al., 2024a). A particularly contentious debate is whether LLMs exhibit Theory of Mind, the ability to model others’ mental states which may be different than their own (Premack and Woodruff, 1978). Several studies have shown evidence for (Kosinski, 2023) and against (Sap et al., 2022; Ullman, 2023) Theory of Mind, including studies that develop more rigorous evaluation methods via procedural generation (Gandhi et al., 2024) and adversarial examples (Shapira et al., 2024). Our analysis provides a quantitative approach to engaging with this debate, as inverse decision-making can be considered a specific form of Theory of Mind.
@@ 2.2 Forward and inverse models of human decision-making
One of the most classic and extensively studied problems in decision-making is the risky choice task (Kahneman and Tversky, 1979; Edwards, 1954; Bourgin et al., 2019), where people choose among gambles with different outcome probabilities and payoffs. The rational action is to choose the gamble with the highest expected value, calculated by summing the product of the probabilities and the values of the outcomes. On this task, humans have been described as deviating from rationality in a fourfold pattern: risk seeking for small probability gains, risk averse for small probability losses, risk averse for moderate or large probability gains, and risk seeking for moderate or large probability losses (Kahneman and Tversky, 1979). Other studies have shown that humans tend to act in accordance with bounded rationality, where rationality is traded-off with mental effort, information, and time (Evans et al., 2015; Alanqary et al., 2021).
Peterson et al. (2021) collected the choices13k dataset, a large human dataset with over 13,000 risky choice problems, and showed that people’s decisions in this setting could be captured by simple machine learning models. Binz and Schulz (2023a) used this dataset to fine-tune an LLM, achieving similar performance. Chen et al. (2023) built a risky choice dataset and found that GPT-3.5 makes economically rational decisions, which we replicate in task 3 of our forward modeling experiments on the choices13k dataset across various LLMs.
People are able to infer an agent’s beliefs, desires, and percepts from their actions (Baker et al., 2017; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021). These inferences are typically modeled by assuming that people employ a forward model of decision-making — typically a noisy rational model — and use Bayes’ rule to invert that model (for more details, see Section 4). The study that we focus on, by Jern et al. (2017), was intended to directly test this assumption. Similar approaches have been used to align AI systems to preferences inferred from observed user interactions (Fränken et al., 2023) or to improve human-AI interaction (e.g., Dragan et al., 2013; Sadigh et al., 2016). More generally, inverse modeling can also be viewed as inductive reasoning from behaviors to utility functions, where LLMs have been shown to be skilled at proposing hypotheses to explain observations, but not at applying these hypotheses to novel examples (Qiu et al., 2024). Lastly, recent work has extended forward and inverse models using LLMs to make inferences from utterances as well as actions (Zhi-Xuan et al., 2024; Ying et al., 2023).
@@ 3 Forward modeling: Predicting which gamble people will choose
Our forward modeling experiments used the risky choice paradigm, one of the most basic and extensively studied problems in psychological decision theory (Kahneman and Tversky, 1979; Edwards, 1954). In this paradigm, participants face a choice between two gambles $A$ and $B$, each with a set of outcomes that differ in their payoffs $\mathbf{R}$ and probabilities $\mathbf{Q}$. For instance, a risky choice problem might ask, “would you rather win \$5 with probability 1 or take a 0.5 probability of winning \$10?”. This can be formalized as gamble $A$ having a single outcome with reward $\mathbf{R}_{A}=[5]$ and probability $\mathbf{Q}_{A}=[1]$, and gamble $B$ having two potential outcomes with rewards $\mathbf{R}_{B}=[10,0]$ and probabilities $\mathbf{Q}_{B}=[0.5,0.5]$. Given a risky choice problem, the goal is to find a probability $P(A)$ that is consistent with how likely people would select gamble $A$ over gamble $B$.
To understand how LLMs empirically capture human intent and align with actual human decisions, we designed three forward modeling tasks. First, we asked LLMs to predict the decisions that a human participant would make. Second, we asked LLMs to predict the proportion of participants that would select each option. Third, we instructed LLMs to simulate participants by making the decisions themselves. The second task corresponds directly to the original objective of finding probability $P(A)$ consistent with how likely people would select gamble $A$ over gamble $B$. In the first and third tasks the LLM outputs a binary decision $\{A,B\}$, which is repeated many times to compute an aggregate probability estimate of choosing gamble $A$, which is compared against $P(A)$.
@@
Human choice data came from the choices13k dataset (Peterson et al., 2021; Bourgin et al., 2019), a comprehensive collection of 13,006 risky choice problems. Each choice problem was answered by 15 or more participants, and participants answered each choice problem five times. For each problem, the dataset included the proportion of participant answers that selected each option. Our analyses used a subset of 9,831 problems that were not “ambiguous” (where probabilities were not shown) or lacked feedback. We used the data to evaluate the alignment of LLMs with actual human choices in each of the three paradigms.
@@
We evaluated the following open-sourced and closed-sourced models: Llama-3-8B, Llama-3-70B, Claude 3 Opus, GPT-4-Turbo (0125-preview), and GPT-4o. We implemented experiments on the full choices13k dataset for zero-shot and CoT prompting across the three different tasks mentioned above for all models besides Claude. For Claude 3 Opus, we only evaluated the first task — predicting what a human participant’s choice might be — due to cost limitations.
For the experiments predicting and simulating individual human decisions (tasks 1 & 3) and models Llama-3-8B, Llama-3-70B, and GPT-4-Turbo, we conducted zero-shot experiments with $n$ completions, where $n$ is the number of participants that made the same decision in the choices13k dataset (ranging from 15 to 33). Each completion involved predicting or simulating a single human participant’s decision, with temperature set to 0.7 to maintain sample diversity. For the corresponding CoT experiments, we observed that responses were completely deterministic at temperature 0.7 on a random subset of 1000 decisions, and thus prompted for 1 completion with temperature 0.0. For the same reason, we ran the experiments predicting the proportion of humans who would select an option (task 2) for 1 completion for both zero-shot prompting and CoT prompting with temperature 0.0. To validate our findings, we also ran ablations varying temperature, which we report in Appendix E.1.
We adapted a prompt template previously used by Binz and Schulz (2023a, see Appendix A for example prompts). For each choice problem, we first introduced the decision context of the choices13k dataset (i.e., the idea of choosing between gambling options) before providing each option’s probabilities and associated rewards in dollars. Finally, we asked what the participant(s) would choose (for predicting decisions) or what “you” would choose (when simulating humans). To minimize any positional biases (Wang et al., 2023), we shuffled the order of the options presented in the prompt. We also conducted an ablation where we assigned demographic profiles to the decision maker to show that our results are consistent across prompt variations (see Appendix E.2).